OpenStudy (anonymous):

Which of the following has a slope of –2 and a y-intercept of 4?

OpenStudy (anonymous):

y = –2x – 4 y = 2x – 4 y = 2x + 4 y = –2x + 4

OpenStudy (anonymous):

I love these sort of questions! The equation is y=mx+b m=slope(in this case -2) and B is your y intercept (In this case 4) Can You fill out the equation using this information now?

OpenStudy (anonymous):

Let me write it out a bit better \[Y=mX+b\]